You have two summer jobs. You mow lawns for $20.00 per hour. You also work at a restaurant for 7.50 per hour. In one week you earn 825.00 working 50 hours. Write an equation that can be used to find the hours that you mowed lawns.
Answer by stanbon(75887)

You have two summer jobs. You mow lawns for $20.00 per hour. You also work at a restaurant for 7.50 per hour. In one week you earn 825.00 working 50 hours. Write an equation that can be used to find the hours that you mowed lawns.
------------------
Equations:
L + R = 50
20L + 7.50R = 825 dollars
-----------------------------------
Substitute to get an equation in "L".
---
Since R = 50-L you get:
20L + 7.50(50-L) = 825
-----------------------------
Cheers,
Stan H.
PS: Reply to stanbon@comcast.net